5. (3 points) The General Social Survey asked 1000 people how many hours per day la- they were able to relax. Consider these 1000 people to be a population, and let distribution for the number of hours of relaxation time r. Number of Hours1 a be the number of hours of relaxation for a person at random from this popu tion. The results are presented in the following table. Complete the probability 6 8 3 Frequency | 54 | 96 | 166 245 175 135 85 54 Probability 1000 1000 of 044 Relaxation Time .054 (a) (2 points) Find the probability that a person relaxes at most 3 hours. (b) (2 points) Find the probability that a person relaxes at least 3 hours. (c) (2 points) Find the probability that a person relaxes from 3 to 5 hours. (d) (2 points) Find (e) (2 points) Find ơ
Answer #1

5.

First of all we complete the table :

Number of hours 1 2 3 4 5 6 7 8
Frequency 54 96 166 245 175 135 85 44
Probability of Relaxation time 0.054 0.096 0.166 0.245 0.175 0.135 0.085 0.044

Let X be the Number of hours of Relaxation time.

(a)

Probability that a person relaxes at most 3 hours = P(X=1) + P(X=2) + P(X=3)

= 0.054 + 0.096 + 0.166 = 0.316

(b)

Probability that a person relaxes at least 3 hours = P(X=3) + P(X=4) + P(X=5) + P(X=6) + P(X=7) + P(X=8)

= 0.166 + 0.245 + 0.175 + 0.135 + 0.085 + 0.044 = 0.85

(c)

Probability that a person relaxes from 3 to 5 hours = P(X=3) + P(X=4) + P(X=5)

= 0.166 + 0.245 + 0.175 = 0.586

(d)

に8 μ = Σ!P(X = z) = 1 * 0.054 +2 * 0.09G+ + 8 * 0.044 4.356

(e)

sigma = sqrt{ sum_{x=1}^{x=8}x^2P(X=x) - left (sum_{x=1}^{x=8}xP(X=x) ight )^2}

Στ2PlX = z) = 12 * 0.051+ 22 * 0.096 +8 0.044-22.068 に!

and we already calculated :

Σ1P(X = z) = 1 * 0.054 +2 * 0.096 + + 8 * 0.044 4.356 に!

ơ 、/ 22.068-(4.356)2-1.7